OpenStudy (anonymous):

2. A solution contains 85.0 g of NaNO3, and has a volume of 750 mL. Find the molarity of the solution. You MUST show all work to receive full credit. (3 points)

OpenStudy (anonymous):

M=mass/m.mass X 1000/V in ml using this formula calculate it

OpenStudy (anonymous):

molar mass of NaNO3=96 g/mol mass is = 85 g Vol in ml = 750 ml M=85/96 X 1000/750=1.18 molar

OpenStudy (anonymous):

yes